# Visualization of Electronic Health Record Sequences at Scale

### v1

@@ 3.4 Optimization of ESS Visualization Programs
Table 1: CSV file showing records using our format, typical of event sequence systems. Patient #0 is 71 and bought the drug ‘‘Dutastéride’’ containing the 5-alpha molecule in May and June 2015. He suffers from diabetes.
@@
23 current_node = tree # start from the root
24 context = Context() # we need to maintain a context for aggregating events
25 for event in lowlevel+ENDEVENT: # Add a fake empty event at the end
26 for highlevel in context.low_level_to_high_level(event):
27 # low_level_to_high_level returns 0 or more events
28 # highlevel looks like: {type: "alphabloc", age: 42, duration: 300}
29 current_node = merge_node(current_node, highlevel)
Listing 1: Overview of a non-progressive ESS visualization.
[…]
The first optimization relies on the memory layout, which organizes the low-level event sequences in consecutive memory positions so the aggregation algorithm can access them sequentially. Modern hardware strongly optimizes this access pattern by prefetching memory caches along with sequential access. The aggregation process takes a table of low-level events indexed per patient and outputs a prefix tree meant to be rendered. Note that the run time is essentially linear in the number of events, but existing implementations use data structures that are expensive both in memory and execution time, limiting the dataset size allowed to remain interactive. In EventFlow, transforming low-level events into high-level ones is expensive due to the rich set of possible transformations supported, as shown in the Supplementary Material. We write the function low_level_to_high_level directly in C++ for two reasons: expressive power and optimization. This function is specific to a treatment or use-case (e.g., non-cancerous prostate adenoma, used as an example here), so we hard-code the rules, e.g., the definition of what is an “interruption” synthetic event in a specific treatment. We thus make sure the rules follow state-of-the-art recommendations. We allow the analysts to change some parameters within meaningful limits during the exploration. The rules are more complex than what EventFlow can express, yet rather simple to express using a regular programming language like C++. ParcoursVis’s implementation is 190 lines long. Adapting it to another scenario, described in Section 5, took a few days and had a similar length.
[…]
Second, for the progressive aspect, we rely on the patients being shuffled. Shuffling is important in PVA for faster convergence and to overcome order bias in the data [39]. Keeping patients in chronological order would initially reveal pathways that followed older medical protocols and delay the visualization of more recent ones, biasing the early overview snapshots.
The main loop performs the tree aggregation along with the progressive calculations of the distributions. As shown in LABEL:lst:aggregation line 6, the back end partially aggregates the tree with a certain number of sequences (a chunk) and sends it to the front end (function render in LABEL:lst:aggregation line 8), and when the function returns (the user interface received the data), it continues the aggregation for the next chunk.
To determine the chunk_size parameter for a given iteration (kept constant in the pseudo-code of LABEL:lst:aggregation), we measure the average speed (patients per second) of the last six iterations and multiply it by the desired latency (typically 1 second). We start with the pessimistic chunk size of 100,000 patients and converge to a chunk size adapted to the network latency and performance of the back end.
We send the resulting tree to the rendering client in JSON format after two transformations: the duration lists are transformed into histograms, and small nodes below a specified threshold of patients (10–50) are trimmed. In our applications, the final tree has around 2k nodes; it can be transferred in a few milliseconds to the front end through a web socket.
[…]
Rendering is done on the front end in two passes: first, it preprocesses the tree according to user-specified parameters, such as the maximum tree depth and event filters, generating a rendering tree. Then, it lays out the Icicle tree in SVG format that eventually appears on the screen. The first step rewrites the aggregated tree into a new one when only simple transforms are involved, saving back-end heavy work. Typically, to filter out node types or apply a hierarchy transform (see Section 3.2.2 for details). This allows for a smoother user experience since a tree is usually composed of a few thousand nodes that are fast to process compared to the dataset that can contain millions of events.
To prevent the progressive algorithm from flooding the rendering, the front end notifies the back end to continue the aggregation process once it has fully displayed an updated tree. When the user changes any of the parameters used for aggregation, the aggregation algorithm restarts from the beginning, iteratively updating the rendering.

### v2

@@ 2.1. Temporal Event Sequences
Table 1. CSV file showing records using our format, typical of event sequence systems. Patient #0 is 71 and bought the drug ‘‘Dutastéride’’ containing the 5-alpha molecule in May and June 2015. He suffers from diabetes.
Our tool, ParcoursVis, aims to support domain experts in visualizing and exploring patients’ care pathways. Like EventFlow (Monroe et al., 2013), it uses aggregation and filtering to visualize event sequences such as the ones shown on Table 1, extracted from the SNDS (SNDS, 2023). According to the taxonomy in Wang et al.’s survey of EHR visualization (Wang and Laramee, 2022), it falls into the “Event Sequence Simplification” (ESS) family of techniques (explained in Figure 2). This family is defined as “any technique used for reducing the visual complexity of event sequences in aggregated display overviews” (Wang and Laramee, 2022). Wang et al. distinguish seven families: ML, NLP, ESS, Geospatial Visualization, Clustering, Comparison, and Others. Among the 51 articles related to the visual analysis of EHR data, 11 belong to the ESS family. In other families, visualizing an overview of the event sequences is not the focus; they rather focus on one patient, a set of patients, or spatial areas sharing some characteristics.
In addition to visualization, EHR data is also analyzed using sequence mining augmented with visualization (Vrotsou and Nordman, 2019; Liu et al., 2017; Perer and Wang, 2014a; Stolper et al., 2014). The idea consists of trying to automatically discover frequent and important patterns in the sequences to discover unexpected information. This approach differs from ours since it only needs to visualize the search results. Instead, our tool interactively visualizes all the treatment pathways following Shneiderman’s mantra: overview first, zoom, filter, and details on demand (Shneiderman, 1996).
$a_{2}$$b_{2}$$c_{1}$$d_{1}$$e_{1}$
$a_{2}$$b_{2}$$c_{1}$$d_{1}$$e_{1}$
Figure 2. ESS Aggregation of EHR Sequences as implemented by EventFlow and ParcoursVis. Patient A took the treatment sequence $\{a,b,c\}$ in that order, and patient B took $\{a,b,d,e\}$. The left side shows the prefix tree of these sequences, and the right side shows the Icicle tree visualization where node height encodes the frequency. Node width is constant here, but, in EventFlow and ParcoursVis, it encodes the mean duration of the treatments.:
2.1.1. Aggregation of EHR Data
Figure 2 explains how a list of sequence of events is aggregated into a tree. In addition, there is also an aggregation at the sequence level, applied by both EventFlow and ParcoursVis for creating high-level events, which are the ones used to build the prefix tree.
An EHR database, like the SNDS (SNDS, 2023), contains low-level events associated with each patient. For example, a patient treated for hypertension needs to buy a box of pills containing a specific molecule every three months. Low-level events in the SNDS database (SNDS, 2023) include, for instance, the drug’s date of purchase and the associated molecule ( Table 1). When the patient purchases the same box repeatedly over time (it can be up to many years), as shown for patients 0 and 1 on Table 1, we transform the sequence into one high-level event (refered simply as an event) with a start-date and end-date (i.e., a duration) associated with the active molecule, whatever the brand. Here, patient 0 experiences the event “fivealpha”, with a duration of 6 days (from 2015-05-29 to 2015-06-05). This kind of aggregation rule simplifies massively the prefix tree and is standard in EventFlow and ParcoursVis.
ParcoursVis can handle more complex aggregation rules for mixing drugs or generating synthetic events when treatments are interrupted or abandoned. For example, in the case of prostate adenoma treatment, shown on Figure 1, patients can take different types of molecules at the same time, such as alphabloc and fivealpha. The visualization builds and shows the event “fivealpha alphabloc” accordingly. In other contexts, two consecutive low-level events with different molecules are interpreted as switching from one treatment to another. We also create synthetic events to track adherence to treatment when a patient takes too long before buying the next box of treatment (“interruption” event), or stops their medication (“no more treatment” event). When setting up the visualization for a particular pathology or use case, an expert needs to extract the relevant low-level events and specify the different rules to apply.
Given the large number of events available in databases such as the SNDS (SNDS, 2023), which hosts the medical records of the entire French population, designing their visualization to remain interactive at scale is a challenge. Next, we introduce what scalability is and how to characterize it, and review the state of scalability in related applications.
2.2. Scalability
2.2.1. Progressive Visual Analytics
PVA is the key technique we use to guarantee that ParcoursVis is interactive for very large dataset sizes. PVA allows users to visualize the result of an algorithm while it is computed, in a progressive way. Instead of processing the whole data at once, waiting an unpredictable time, as standard algorithms do, PVA algorithms process data chunk by chunk (or iteration by iteration in some cases) (Fekete et al., 2024) to control the latency and show intermediate results iteratively at a controlled pace. The main goal is to keep the overall system interactive and yield results within the typical user’s attention span that lasts 1–10 seconds (Nielsen, 1993; Zgraggen et al., 2017). Practitioners already applied and studied the PVA paradigm on multiple types of visualizations, as described in the survey of Ulmer et al. (Ulmer et al., 2024).
In this work, we progressively aggregate event sequences of patients’ care pathways into a prefix tree (Wang and Laramee, 2022) that we visualize as an Icicle tree. Although PVA has been used for pattern mining in EHRs (Stolper et al., 2014), to the best of our knowledge, we are the first to apply PVA to visualize EHR aggregated data at scale. We also introduce the first progressive tree visualization technique, according to the survey of Ulmer et al.
By quickly presenting partial results, PVA systems allow users to get quick feedback on the overview and queries; it separates latency from completion time. It was also shown to increase users’ commitment and actions per minute (Zgraggen et al., 2017). In the next subsection, we describe the scalability characterization introduced by Richer et al. (Richer et al., 2024b). Based on their definition, we show, in this article, that our PVA algorithm for computing the prefix tree is indeed scalable: the aggregation time scales linearly with the number of events, inversely proportional to the number of cores, and updates always meet the requested latency.
2.2.2. Modeling Scalability
Richer et al. (Richer et al., 2024b) characterize scalability in visualization using a conceptual model that expresses the scalability of a visualization system by a function $f:(S;R,A)\longmapsto E$ with four components: the problem size $S$ (i.e., set of variables attuned to the size of the input data), resources $R$ (i.e., hardware or application environment used for the visualization), assumptions $A$ (i.e., boundaries of the research context and problem definition), and effort $E$ (i.e., measures of the claimed system’s performance).
In our case, $S$ relates to the number of patients and the number of events per patient. The two numbers are proportional if we consider patients to have an average number of events, but depending on the pathology studied, the average can vary from ten (e.g., for emergency service sequences) to thousands (e.g., for lifelong chronic pathologies). $R$ is related to the machine we use and the number of CPU cores used by the computation. Assumptions $A$ appertain to (1) the data, (2) the technical setup, and (3) the preprocessing.
  1. (1)
    We limited the number of event-types (e.g., ten in Figure 6, between eight and twenty in other applications) since event types are encoded with color, and maintaining a relatively small set ensures that the primary events remain sufficiently distinguishable. Note that the event type distribution is always skewed, as shown in Figure 6.
  2. (2)
    We use a web-based application where a back-end machine serves the aggregated data through the internet, featuring a network bandwidth that can be low due to hospital wifi congestion, and a front-end local browser to interact with the visualization. We exclude the use of a GPU in the browser since we do not control the front end machines.
  3. (3)
    We also perform some preprocessing ahead of time in the back-end, without precomputing all the possible interactive queries, as our datasets can be filtered on a large combination of attributes. As for $E$, we consider three efforts: latency, completion time, and stability. Note that in non-progressive applications, latency is equivalent to completion time and that stability is specific to PVA and streaming visualizations (Ulmer et al., 2024) (see Section 2.2.1 and Section 2.3).
The scalability model of Richer et al. also introduces multiple ways of characterizing the scalability they call expression and meaning, such as asymptotic scalability using the $O(\cdot)$ notation, or simply claiming to be faster than another system. In our case, Latency should always be under a specified limit to guarantee that users do not wait during the progressive updates of the system, and completion time should be minimized. Also, during the progressive updates, we want to maximize the stability of the visualization to allow users to scan the Icicle tree without sudden changes. We measure the first two efforts by varying the size $S$ of a test dataset and the number of CPU cores $C$, obtaining a scalability shape, i.e., a graph $T=f(S,C)$. The latency between two updates should remain under the 2 s boundary for all problem sizes and resources ($S$ and $C$). Since the algorithm needs to process all the events, the completion time cannot be better than linear with $S$ and $C$, and short. As for the stability, it can be measured in terms of variations or accidents per unit of time, but also as feedback from users because, ultimately, the stability might be considered good enough if the users do not notice instability. We report on both measures.
2.2.3. Scalability in EHR Visualization
Study Reported size Unit Scalability claims
Bernard et al. (Bernard et al., 2015) 65 patients Unspecified
Klemm et al. (Klemm et al., 2015) 1,186 subjects Unspecified
Perer et al. (Perer and Wang, 2014b) 2,336 patients Unspecified
Gotz et al. (Gotz and Stavropoulos, 2014) $2,899$ patients Unspecified
Guo et al. (Guo et al., 2018) 5,800 patients Unspecified
Monroe et al. (Monroe et al., 2013) $\approx 10,000$ patients Latency (same as Time)
Jiang et al. (Jiang et al., 2016) 833,710 cases Unspecified
ParcoursVis >20,000,000>20,000,000 patients Latency, Time, Stability
Vrotsou & Nordman (Vrotsou and Nordman, 2019) 6,477 sequences Unspecified
989,925 sequences Unspecified
Liu et al. (Liu et al., 2016) $\approx 100,000$ sequences Unspecified
Table 2. Reported dataset sizes and scalability claims for the visualization of EHR and other sequence data.
Existing systems that visualize event sequences, both in medical and non-medical domains, often fail to report their scalability relative to the dataset sizes they handle (Richer et al., 2024b). Wang et al.’s survey (Wang and Laramee, 2022) does not report dataset sizes, although it reports scalability as the first future research challenge. In the medical domain, reported dataset sizes vary widely, as shown in Table 2. EventFlow handles datasets with a limited number of sequences (Bjarnadóttir et al., 2016; Monroe et al., 2013; Ozkaynak et al., 2015; Du et al., 2017); about 20 k and with a few seconds latency according to our experiments (available in the Supplemental Material). For handling larger datasets, its manual suggests the strategy of random sampling sequences (Du et al., 2017). Random sampling, however, loses rare and sometimes important sequences that interest practitioners. With ParcoursVis, we found sequences concerning 3 k patients in our multi-million patient datasets; they are relatively rare, but not considered outliers by doctors. Instead, to be scalable and accurate, we use PVA. Some other systems, such as DecisionFlow (Gotz and Stavropoulos, 2014), mention handling over a million individual point events (2,899 patients). However, they provide no details on their scalability. In non-medical domains (mostly website visit sequences), studies report dataset sizes in the range of thousands to millions of sequences, but provide no scalability measure. Overall, for systems dealing with event sequence data, there is either a lack of information on performance relative to dataset size or a focus on small datasets, making it difficult to assess scalability and leaving open the question of how such systems can handle large-scale, real-world event sequence data. To exemplify the scalability of our approach, we visualized the largest dataset mentioned with ParcoursVis (see the .
In contrast to event sequences, tabular data visualization systems provide well-documented scalability benchmarks. They typically rely on online analytical processing (OLAP) over one or several dimensions using aggregation operators such as mean, variance, sum, or count. Some systems, such as imMens (Liu et al., 2013) (which relies on GPU acceleration) and Mosaic (Heer and Moritz, 2024) (which relies on the optimized analytical database DuckDB (Raasveldt and Mühleisen, 2019)), scale to billions of rows. However, event sequence aggregation relies on a different data model and more complex aggregation rules, as explained in Section 2.1.1, rather than these generic operators, making OLAP techniques incompatible with our data. Consequently, their scalability claims are not directly comparable to those for event sequence visualization systems.
2.3. Uncertainty and Stability
PVA systems start by showing partial results, implying uncertainty due to the progressive aspect of the computation. This progressive uncertainty (Vilanova et al., 2024) is reduced over time until the final result is computed, without any progressive uncertainty but potentially with other types of uncertainty, just like standard data and visualizations. The visualization shown early may or may not be close to the final result. It is, therefore, important for analysts to assess the quality of the progressive results to decide if they can trust the visualization or if they should wait longer. Angelini et al. (Angelini et al., 2018) explain that progressive visualizations undergo three stages: (1) early partial results that are usually noisy and not trustworthy, (2) mature partial results that are stabilizing and can be trusted but with uncertainty, and (3) definitive partial results that are accurate and can take (wasted) time to finish. When our aggregation tree is computed progressively, all the nodes will undergo these three stages. Yet, the nodes closest to the root will become reliable and stable earlier than the deeper nodes since they are aggregating more sequences, and that stability is essentially related to the number of sequences they represent.
The uncertainty can be computed and visualized explicitly (e.g., through error bars), or implicitly through measures such as stability. Instability at a given stage indicates that the result is still uncertain, though the reverse is not always true: a progressive visualization may remain stable for some time and still change later (Patil et al., 2023; Correll and Gleicher, 2014). However, explicitely conveying uncertainty remains challenging, even for statisticians who often struggle to interpret it correclty in static visualizations (Correll and Gleicher, 2014). For progressive visualization, many representations are possible, e.g., bar charts with error bars, violin plots, or gradient distribution (Correll and Gleicher, 2014; Fisher et al., 2012), but few have been studied and validated empirically. Patil et al. (Patil et al., 2023), for instance, evaluated four designs for visualizing uncertainty on progressive bar charts and found two to be effective. However, there is no study about the progressive visualization of an Icicle tree, to the best of our knowledge.
ParcoursVis takes less than 20 s to update its tree fully for our 10 M patients dataset, with an update every 2 s. It seems unlikely that users could make sense of any uncertainty visualization added to many nodes in such a short update time; it might take more time to make sense of the uncertainty visualization than to wait for the progressive visualization to complete. Instead, we rely on stability as a proxy for quality, assuming that subtrees will remain stable when the progressive uncertainty decreases (Vilanova et al., 2024). On the other side, when a subtree is not stabilized for a few seconds, the user will not be able to infer much from watching it. We study the convergence and stability of ParcoursVis’s progressive Icicle tree visualization, instability reflecting progressive uncertainty.
3. ParcoursVis
ParcoursVis is a web-based application inspired by EventFlow that processes data in a back end; users interact via a web browser front end. While ParcoursVis is specialized for exploring and analyzing EHR sequences, its core structure is more general and can handle the same kinds of event sequences and applications as EventFlow. Contrary to EventFlow, the rules for aggregating and merging events are written in C++ for performance and expressive power reasons. For real pathologies, many constants and aggregation rules need to be set according to state-of-the-art recommendations that cannot be expressed within EventFlow. Therefore, an engineer is needed to configure ParcoursVis to define new domain-specific rules when it is applied to a new medical question with specific kinds of events.
3.1. Example Use Case
ParcoursVis is currently used in two main applications: (1) non-cancerous prostate adenoma treatment analysis for the French social security, and (2) analysis of patient pathways across all public Parisian hospitals’ Emergency Departments (EDs). While the detailed examples in this paper come from the first use case, we briefly describe the second.
(1) Specific to our non-cancerous prostate adenoma use case, our application receives information from the Social Security reimbursement database (available in the SNDS (SNDS, 2023)) about the drugs bought by patients (time, name, molecule), dedicated surgery, and deaths. As explained in Section 2.1.1, there are three molecules possibly administered for treatment: alpha-blockers, specific herbal medicine, and 5-alpha, respectively labeled alphabloc, phyto, and fivealpha in Figure 1. Those are the three low-level events our referring doctors selected to extract for their analysis. They also extract prostate surgery events and deaths. These events are aggregated into high-level events with durations that are transformed into an Icicle tree and visualized. Our dataset also contains patient’s attributes such as their age and information related to known comorbidities (e.g., diabetes, hypertension) to explore possible correlations with treatment pathways.
(2) We also use ParcourVis for visualizing patient pathways in the 16 public Emergency Departments in Paris. For over 15 years, EDs have faced a steady increase in patient volume, leading to frequent overcrowding and prolonged waiting times. Analyzing patient care pathways offers the potential to identify specific stages where the process becomes delayed, resulting in extended waiting times or medical accidents. Our project deals with 1.5 M patient visits over three years. To use the secure JupyterLab environment provided by the Paris Public Hospitals, we wrapped the ParcoursVis front end into a widget to provide an easy-to-access interface for our collaborators. Emergency Department staff use a centralized system to document patients’ medical records upon their arrival. Physicians enter administrative information (e.g., name, gender, age) and medical procedures (e.g., triage score assessment, medical exams). They also document patient location within the hospital as patients are redirected to different stages of care (e.g., to a standard cubicle or waiting areas). This system saves data every 15 minutes, providing near-real-time location data on each patient throughout their stay. As in the prostate adenoma use case, low-level events (here, changes in patient location within the ED, recorded every 15 minutes) are aggregated into higher-level stages of care, such as, for instance, being located in a waiting area (where patients remain before consultations or results), the triage room (where priority scores are assigned) and the shock room (Intensive Care Unit). In addition to these locations, we include attributes such as patients’ age, the name of the hospital they visited, and the triage score they have been given. The triage score, ranging from 1 (high priority, life-threatening) to 5 (low priority, non-urgent), is assigned to patients based on symptoms, medical history, vital signs, and clinical observations.
Tasks ParcoursVis EventFlow
Simple Tasks Overview: visualize all the temporal event sequences (high-level event types). æ æ
Browse explicit sequences: enable record-level inspection and drill down into individual records. œ æ
Zoom and Filter Events: visualize all the temporal sequences having a given prefix (children of a selected node) or filter out specific event types. æ æ
Details on demand: visualize all the attributes and distributions related to a specific sequence (e.g., age distribution and duration distribution of the patients reaching a node). æ æ
Advanced Tasks System Configuration: the ability to configure the system to perform different aggregations or data preparation. œ æ
Filter by attribute: simplify the visualization by filtering the data based on the attributes and metadata of the sequence (e.g., total duration, age of the patient), and the resulting tree (e.g., minimum size of the nodes). æ æ
Configure view: setting an alignment prefix to visualize what had happened before and after a specific sequence. æ æ
Abstracting: regrouping multiple types of events into a super-type, using a hierarchy. æ æ
History and comparison: navigate through the history, i.e., previous filters, zooms, or alignments. Compare different trees corresponding to different history states by switching between them. æ œ
Temporal queries: provide a temporal query language to search for e.g., the absence of treatment or nested treatments. œ æ
Analytical Tasks Application questions: identify trends, discover frequent, critical, or surprising pathways, and understand underlying causes, finding actionable strategies for improvement or resolution. æ æ
Data questions: identify data errors, limitations, and determine more suitable data extraction methods. æ æ
Table 3. A summary of the different tasks handled by ParcoursVis, outlining the differences with EventFlow.
3.2. User Interface and Tasks
Figure 3. The “Sequence” view visualization (squared) shows all events leading to a specified prefix, here the “surgery” event. About 60% of the patients who underwent surgery were treated before. A button allows showing all events that came after the prefix instead. :
Our visualizations and user interfaces (GUI) are close to EventFlow, which has already been validated through many applications and case studies (Bjarnadóttir et al., 2016; Monroe et al., 2013; Ozkaynak et al., 2015). We summarize the different tasks supported in Table 3. In the remaining, we focus on the specific features introduced by PVA.
3.2.1. Main View and Prefix View
The “Main” view (Figure 1) shows the aggregated tree of all patients’ care pathways. The node width, height, and sorting order are explained in the Figure caption. The color of a node encodes its high-level event type. Clicking on a node shows details on demand in the “Detail” tab of the control panel (4(a)).
Users can change the view and focus on a specific event or sequence of events. The “Prefix” view (Figure 3) shows the sequence tree that leads to or follows a given prefix, specified interactively. Each of these views is updated progressively for scalability to guarantee interactive latency.
3.2.2. Control Panel
(a) “Detail” Tab showing histograms and bar charts for the selected node.
{subcaptiongroup}
Figure 4. “Filtering” Tab showing widget to set parameters and filter on attributes.
Figure 5. “History” Tab showing the history tree for navigation and comparison.
Figure 6. Control Panel Tabs of ParcoursVis. See Figure 1 for the Overview Tab.:
Overview:
The “Overview” tab (Figure 1) shows (1) the number of high-level events aggregated and (2) the hierarchy associated with the data using an Icicle tree. Hovering a node shows the number of patients it contains. Those numbers are updated continuously when the progressive algorithm is running. Clicking on an event hides it from the tree. Clicking a second time makes it appear again. These operations are done in the front end and are instantaneous.
The hierarchy groups multiple high-level events into a higher category, e.g., all drugs under the category “medication”. Figure 1 visualizes the hierarchy and its distribution in the partial aggregated tree. For applications with many event types, using higher hierarchy levels simplifies the visualization. Users can modify the level of hierarchy at any time; the change is instantaneous, as it only impacts the prefix tree visualization; it does not need to restart the progressive aggregation.
Detail:
The “Detail” tab (4(a)) shows the data distributions of a selected node. Currently, we show the age and duration distributions using histograms and the comorbidities using a bar chart. Those graphs are updated iteratively when the progressive algorithm runs.
Filtering:
The “Filtering” tab (Figure 6) allows selecting a subset of patients and filtering the view according to parameters and node attributes. When users adjust the aggregating parameters (e.g., the thresholds defining the interruption events) or select a subset of patients by specifying query parameters, the progressive algorithm restarts from the beginning, considering only sequences matching the query and updating the rules that aggregate low-level events into high-level events. In contrast, filtering the view only manipulates nodes from the already computed tree (e.g., hide small nodes or merge nodes into higher hierarchical levels); it does not process the data again and results in instantaneous updates. To filter the view, users can specify a hierarchical level, truncate the tree to a maximum depth (e.g., show only the first three depth levels), and select a minimum number of patients per sequence to hide small nodes, simplifying the tree.
History and Comparison:
The “History” tab (Figure 6) is inspired by the history comparison feature of Pister et al. (Pister et al., 2023) that relies on the Trrack (Cutler et al., 2020) history management library. It shows the history of the user’s actions as a tree; the user can jump to any past state of the exploration by selecting a history node.
The user can compare the current state with another one using the “compare to” button. The user can visualize the aggregated trees from both states in a back-and-forth manner using the “Next” and “Previous” buttons (not shown on the screenshots). The node associated with the current state in the “History” tab has a blue halo, while the node compared has a green halo. Selecting this second state initially restarts the progressive aggregation using the state’s associated parameters. As the state being compared may not have finished processing all patients when switching back and forth between the compared states, parcoursVis initializes a second progressive context for the second state. Thus, it does not restart the aggregation from the beginning, but resumes it from where it was left previously. This avoids unnecessary waiting for the user to get back to where the context was previously and avoids breaking the user’s mental model related to the context because the progressive algorithm may never get back to the exact previous view.
Overall, the PVA aspect of ParcoursVis required us to consider the usability issues of every feature that either (1) forces the aggregation algorithm to restart, (2) conversely, avoids restarting the algorithm when the front end can handle it alone, and (3) affects the user experience in an unexpected way (e.g., maintaining internally two prefix trees updated progressively for the comparison feature in the ”History“ tab).
3.3. Implementation
CSV file
Pre-processingBinary FilePre-processing of the dataset. Performed only once.MemoryMapPVA backendHard-codedaggregating rulesUIViewFront endPVA pipelineFilter TreeFilter DataUpdate TreeDetails on Demand
Figure 7. Architecture of ParcoursVis, organized in a pre-processing stage, a back end, and a front end. The progressive management of the tree is depicted in red.:
Figure 7 describes the architecture of ParcoursVis. Our front-end application is made of about 3,000 lines of JavaScript that uses D3.js (Bostock et al., 2011) and Vue.js (Vue, 2023). This allows us to show visualizations on a regular web browser accessible to anyone and, in particular, to referring doctors without any installation. However, web technologies are not designed for heavy computation. We, therefore, rely on an optimized back-end library written in about 3,000 lines of C++ and used as a Python module that progressively aggregates the data sent to the front end. The back-end Python server code is 127 lines long. We also provide test suites for JavaScript and C++ that double the source code size. The code is open-source at https://gitlab.inria.fr/aviz/parcoursvis/.
3.4. Optimization of ESS Visualization Programs
[…]
@@ 3.4. Optimization of ESS Visualization Programs
23 current_node = tree
24 context = Context()
25 for event in lowlevel+ENDEVENT:
26 for hl in context.ll_to_hl(event):
27 # ll_to_hl returns 0 or more events
28 # highlevel looks like:
29 # {type: "alphabloc", age: 42, duration: 300}
30 current_node = add_event(current_node, highlevel)
Listing 1 Simple ESS Aggregation Algorithm.
3.4.1. Optimizing Data Preprocessing
First, as shown in Figure 7, ParcoursVis relies on pre-parsed data stored in binary format, avoiding the expensive step of parsing CSV data at each program start while ensuring all events per patient are sorted chronologically and that patients are shuffled to accelerate the aggregation convergence. This last step is useless in non-progressive systems but essential to optimize convergence and stability in PVA ȷ (Fekete et al., 2024).
Running the preprocessing step takes three minutes on our back end for 50 million events stored in CSV format. It is done once per dataset and generates the binary files that will be loaded when ParcoursVis’s back end starts. When the non-progressive program loads the CSV files in a time linear to the file size (Listing 1 line 2), the progressive back end loads the binary files in negligible time using memory mapping (Listing 2 lines 3–4). It is an operating system mechanism that places a local file into main memory instantly from the running program’s perspective.